OpenStudy (anonymous):

how do you find d^2y/dx^2 of x^2+y^2=1

jimthompson5910 (jim_thompson5910):

x^2+y^2=1 d/dx(x^2+y^2)=d/dx(1) 2x+2y*y' = 0 2y*y' = -2x y' = -2x/2y y' = -x/y So the first derivative wrt x is y' = -x/y Now derive both sides again y' = -x/y d/dx(y') = d/dx(-x/y) y'' = d/dx(-xy^(-1)) y'' = -y^(-1) + xy^(-2)*y' y'' = -y^(-1) + xy^(-2)*(-x/y) y'' = -1/y - x^2/y^3 y'' = (-y - x^2)/(y^3) y'' = -(x^2+y)/(y^3) So the second derivative wrt x is y'' = -(x^2+y)/(y^3)

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

np